# Mamba-based Deep Learning Approaches for Sleep Staging on a Wireless Multimodal Wearable System without Electroencephalography

### v1

@@ Abstract
We investigate using Mamba-based deep learning approaches for sleep staging on signals from ANNE One (Sibel Health, Evanston, IL), a minimally intrusive dual-sensor wireless wearable system measuring chest electrocardiography (ECG), triaxial accelerometry, and temperature, as well as finger photoplethysmography (PPG) and temperature.
@@ Methods
We obtained wearable sensor recordings from 360 adults undergoing concurrent clinical polysomnography (PSG) at a tertiary care sleep lab. PSG recordings were scored according to AASM criteria. PSG and wearable sensor data were automatically aligned using their ECG channels with manual confirmation by visual inspection. We trained Mamba-based models with both convolutional-recurrent neural network (CRNN) and the recurrent neural network (RNN) architectures on these recordings. Ensembling of model variants with similar architectures was performed.
@@ Results
Our best approach, after ensembling, attains a 3-class (wake, NREM, REM) balanced accuracy of 83.50%, F1 score of 84.16%, Cohen’s of 72.68%, and a MCC score of 72.84%; a 4-class (wake, N1/N2, N3, REM) balanced accuracy of 74.64%, F1 score of 74.56%, Cohen’s of 61.63%, and MCC score of 62.04%; a 5-class (wake, N1, N2, N3, REM) balanced accuracy of 64.30%, F1 score of 66.97%, Cohen’s of 53.23%, MCC score of 54.38%.
@@ Conclusions
Deep learning models can infer major sleep stages from a wearable system without electroencephalography (EEG) and can be successfully applied to data from adults attending a tertiary care sleep clinic.
@@
3Vector Institute, 4University Health Network Toronto, ON, Canada
5Queen’s University Dept. of Electrical & Computer Engineering Kingston, ON, Canada
@@ 1.1 Sleep Staging and Polysomnography
Sleep is essential to physical and mental health. Quantitative evaluation of sleep architecture and physiology is central to the diagnosis of many sleep disorders and plays an important role in the diagnosis of neurodegenerative diseases [1]. Sleep is traditionally and most reliably measured by polysomnography (PSG), in which brain activity, eye movements, heart rate (HR), muscle tone, and breathing patterns are recorded simultaneously using sensors attached to a patient sleeping in a laboratory environment. This is followed by sleep staging, the manual categorization of time periods into different stages of sleep based on a trained sleep technician’s visual inspection of the collected physiological data. The exact sleep staging methodology depends on the specific framework; under the American Academy of Sleep Medicine’s (AASM) [2] conventions, recordings are divided into 30-second windows, or "epochs", with electroencephalography (EEG), electromyography (EMG), and electro-oculography (EOG) used to classify epochs into three broad stages: awake, rapid eye movement (REM) sleep, and non-REM (NREM) sleep, which can be further subcategorized as N1, N2, and N3 sleep. EEG is the principal signal used for sleep staging [2, 3, 4].
@@ 1.2 Wearable Devices and Automated Sleep Staging
The need for specialized personnel, equipment, and laboratory space makes PSGs expensive to conduct and difficult to scale. The cumbersome sensor setup and unfamiliar resting environment may also be uncomfortable for many patients and can result in sleep that is unrepresentative of that of the home environment. Furthermore, the throughput of PSGs is limited by the time-consuming process of monitoring patients throughout the night and then visually labelling sleep epochs.
In response to these challenges, wearable devices have been developed that record some but not all of the signals recorded by conventional PSG. Devices recording EEG have the advantage of directly measuring brain activity but the disadvantage of the discomfort associated with sensors on the head, similar to PSG. On the other hand, EEG-less devices are less obtrusive, and more easily set up for in-home use without a technician, but infer rather than directly measure sleep stage.
There has been considerable interest in pairing wearable devices with automated sleep stage inference in order to minimize the need for specialized technologists for labeling and hence improve scalability. Indeed, even for the analysis of conventional in-labratory PSG, automated sleep staging has been explored as a means of addressing the time-intensiveness of sleep epoch labelling. The focus on pattern recognition and loosely rule-based nature of sleep staging is ideal for machine learning approaches which have employed model architectures like convolutional neural networks (CNNs) [5], recurrent neural networks (RNNs) [6], hybrids of the previous two (e.g. convolutional recurrent neural networks (CRNN)) [7], Transformers [8], and graph neural networks [9]. Deep learning models trained on PSG data, most notably the EEG signal or some combination of EEG, EMG, and EOG, have surpassed inter-scorer agreement levels, and can make predictions orders of magnitude faster than a human evaluator at the cost of sacrificed interpretability [10, 11].
@@ 1.3 Recurrent Neural Networks and Mamba
RNNs are a class of neural networks that take in a sequence of vectors and produce a vector output per input, allowing them to execute a wide breadth of time-series tasks. Prevalent modern RNN architectures include long short term memory (LSTM) [12] and gated recurrent units (GRU) [13], which are known to forget very long-range dependencies and whose architecture cannot take advantage of GPU acceleration, limiting the training speed.
@@
| Male/Female | 181/176 (51/49%) | ||
| Awake [%] | 43.43 | 17.15 | 11.54 95.24 |
| N1 Sleep [%] | 10.67 | 6.13 | 1.05 33.13 |
| N2 Sleep [%] | 27.18 | 10.73 | 0.58 63.59 |
| N3 Sleep [%] | 10.38 | 7.94 | 0.00 54.83 |
| REM Sleep [%] | 8.34 | 5.81 | 0.00 28.80 |
@@
| Body Mass Index (BMI) [kg ] | 27.81 | 5.73 | 10.00 50.40 |
@@ 1.4 Contribution
This paper applies Mamba-based deep learning approaches to the problem of sleep staging using data from the ANNE One [18] sensor. ANNE One, as seen in Figure 1A, is a flexible, minimally-intrusive, and clinical-grade wireless dual-sensor wearable system by Sibel Health composed of a finger module, which measures photoplethysmography (PPG) and temperature (Temp.), and a separate module attached to the chest via an adhesive, which measures electrocardiography (ECG), triaxial accelerometry (Acc.), and temperature. The raw signals of a typical recording are plotted in Figure 1B.
An anterior study centered on ANNE One [19] exists, which primarily emphasized the development of person-specific models which were trained and evaluated on data from the same patient, with a population model trained and evaluated on data from different patients showing much poorer performance. We aim to develop a better population model that allows accurate sleep staging from ANNE One signals recorded from previously unseen patients.
There is a considerable body of work applying machine learning to the problem of sleep stage inference from non-EEG signals. For the most part, these have been trained and tested using the cardiopulmonary sensors of conventional PSG [20, 21, 22, 23, 24, 25, 26], which are typically artifact-free. Relatively few studies have applied these approaches to data from wearable devices, which are in general more prone to poor signal quality and signal loss; consequently, such studies often need to reject epochs of poorer-quality signal prior to model training. Moreover, these works have focused on analysis of data from single sensors, typically ECG or PPG [27, 28]. Finally, models developed for wearable sensors have largely trained on data from relatively young and healthy volunteers, rather than patients presenting to sleep clinics, raising a question of generalizability to clinical populations.
The current study builds on and extends this body of work in several ways. First, we leverage the ANNE One devices’ multi-sensor nature, which may in principle allow for predictions even in settings where one or more sensors are reading low-quality signals. Second, we train and test on data from relatively older patients with a range of sleep disorders (Figure 1C) rather than younger and healthier patients, ensuring applicability to older adults with sleep complaints. Finally, this work broadens the Mamba architecture to EEG-less sleep staging, and aims to answer whether the incorporation of Mamba offers a relative performance increase similar to PSG EEG sleep staging as noticed in [16, 17].
@@ 2.1 Dataset
ANNE wearable sensors were used to collect full-night recordings from 360 adults undergoing concurrent clinical PSG at a tertiary care clinical sleep laboratory in Sunnybrook Health Sciences Centre, Toronto, Canada. PSG recordings were recorded according to AASM practice parameters for in-labratory PSGs using a Grael PSG system (Compumedics, Victoria, Australia). The study was conducted in accordance with the latest version of the Declaration of Helsinki. All subjects provided written informed consent for the full study protocol.
Participants were first instrumented with the ANNE One, whose recordings were started ahead of the set-up and initiation of PSG recordings. In order to mimic what would occur in a home environment, the ANNE signals were not monitored in real time, and no attempt was made to correct sensor-related issues (e.g. poor skin contact) overnight. Upon completion, each PSG recording was manually labelled by one of four technologists according to AASM standards [2]. In addition, the portion of the ANNE recordings preceding the onset of PSG recording was labeled as wake based on direct observation of participants by study staff. PSG and wearable sensor data were automatically aligned by their shared ECG signals using cross-correlation analysis and verified by visual inspection.
Three participants who did not sleep at all were excluded from analysis, leaving a total of 357 patients. The characteristics of the study’s participants are outlined in Figure 1C.
The machine learning task is thus defined from the dataset as multiclass 1D signal segmentation, wherein the ANNE One signals constitute the features, the PSG-annotated sleep stages form the labels, and class predictions are made per point in time, as demonstrated in Figure 2A.
@@ 2.2 Preprocessing
The ECG data were recorded at 512 Hz and preprocessed as follows. A wavelet transform was applied to raw ANNE ECG signals to identify temporal windows containing QRS complexes; within each window, raw ECG signals were searched for local maxima and minima to identify the precise timing of the R-peak. The wavelet-transformed ECG data were subsequently downsampled to 100 Hz for additional analyses. Meanwhile, the Hilbert transform was applied to extract the envelope of each QRS complex, the area under which being taken as a measure of the signal to noise ratio (i.e. signal quality index (SQI)). For each peak, morphological features (e.g. amplitude, sharpness) were extracted and hierarchical density-based spatial clustering of applications with noise (HDBSCAN) [29] served to group all identified peaks into a small number of clusters. Peaks occurring in trains with relatively stable RR intervals in the physiological range of human HR were identified as likely representing physiological QRS complexes. QRS complexes that clustered morphologically with these peaks, irrespective of regularity, were used to compute beat-to-beat heart rate.
@@ 2.3 Model
Two Mamba-based deep neural network architectures were investigated: a convolutional-recurrent neural network (CRNN) and a recurrent neural network (RNN) as summarized in Figures 2AB, 3. The CRNN is a typical sleep stage architecture [21, 11] that emulates the way human evaluators label PSGs: the convolutional component encodes physiological signals within each epoch while the recurrent component captures transitional relationships between sleep stages across epochs. The RNN architecture is a lightweight and parameter-scarce alternative to the CRNN, consisting of less parameters (580K vs. 70K). Since Mamba is designed for pattern recognition across long contexts, it could potentially interact with signals more effectively than the CRNN if the signals are given directly on a finer time-resolution without windowing the data.
The CRNN architecture has three input channels for time-series, frequency, and scalar data (see Figure 3): the first two are 3D inputs, with axes for the number of 30-second windows, the window size (i.e. 30 seconds worth of sequential data), and number of features, while the 2D scalar inputs have axes for the number of windows and number of features. The model comprises:
@@
A recurrent component, consisting of a bidirectional Mamba block that processes the embedding vectors sequentially, followed by an element-wise multi-layer perceptron (MLP) that diminishes the feature-size axis into logits for each class. The bidirectional Mamba block (Appendix A) is the concatenation of two Mamba blocks that process epochs in opposite directions.
@@
A bidirectional Mamba block that processes windows sequentially.
@@
Our previous investigations [30] (Appendix C) have found engineered frequency domain and scalar features to yield significant performance benefits for GRU-based deep learning approaches. The engineered scalar features distribute differently depending on sleep stage, while changes to their frequency domains often designate sleep stage transitions (Appendix B). Thus, empirical selections of these features (Table 1) were introduced alongside the time-series signals for both the CRNN and RNN.
@@
RNN CRNN Time-Series Frequency Scalar ECG* Heart Rate (ECG + PPG) ECG* Age PPG* Heart Rate (ECG) PPG* Sex X* Heart Rate (PPG) X* Mean, ECG SQI Y* Pulse Arrival Time (PAT) Y* Mean, PPG SQI Z* Euclidean Norm Minus One (ENMO) Z* Entropy, PPG Chest Temp. Z-Angle Z-Angle* Entropy, PPG* Limb Temp. Limb-Chest Temperature Difference Skew, PPG* Theta Scale, Heart Rate (ECG + PPG) Phi Kurtosis, ENMO Entropy, Z-Angle

### v2

@@ Abstract
We investigate Mamba-based deep learning approaches for sleep staging on signals from ANNE One (Sibel Health, Evanston, IL), a non-intrusive dual-module wireless wearable system measuring chest electrocardiography (ECG), accelerometry, and temperature, and finger photoplethysmography (PPG) and temperature.
@@ Methods
We obtained wearable sensor recordings from 360 adults undergoing concurrent polysomnography (PSG) at a tertiary care sleep lab. Each PSG recording was manually scored and these annotations served as ground truth labels for training and evaluation of our models. PSG and wearable sensor data were automatically aligned using their ECG channels with manual confirmation by visual inspection. We trained Mamba-based models with convolutional-recurrent neural network (CRNN) and the recurrent neural network (RNN) architectures on these recordings. Ensembling of model variants with similar architectures was performed.
@@ Results
Our best approach, after ensembling, attains a 3-class (wake, non rapid eye movement [NREM] sleep, rapid eye movement [REM] sleep) balanced accuracy of 84.02%, F1 score of 84.23%, Cohen’s of 72.89%, and a Matthews correlation coefficient (MCC) score of 73.00%; a 4-class (wake, NREM stage 1/2 [N1/N2], NREM stage 3 [N3], REM) balanced accuracy of 75.30%, F1 score of 74.10%, Cohen’s of 61.51%, and MCC score of 61.95%; a 5-class (wake, N1, N2, N3, REM) balanced accuracy of 65.11%, F1 score of 66.15%, Cohen’s of 53.23%, MCC score of 54.38%.
@@ Conclusions
Deep learning models can infer major sleep stages from the ANNE One, a wearable system without electroencephalography (EEG), and can be successfully applied to data from adults attending a tertiary care sleep clinic.
@@
3Vector Institute Toronto, ON, Canada
4University Health Network Toronto, ON, Canada
5Beth Israel Deaconess Medical Center Boston, MA, USA
6Rush University Medical Center Chicago, IL, USA
7Queen’s University Dept. of Electrical & Computer Engineering Kingston, ON, Canada
@@ 1.1 Sleep Staging and Polysomnography
Sleep is essential to physical and mental health. Quantitative evaluation of sleep architecture and physiology is central to the diagnosis of many sleep disorders and plays an important role in evaluating individuals with and at risk for neurodegenerative diseases [1]. Sleep is traditionally and most reliably measured by polysomnography (PSG), in which brain activity, eye movements, heart rate (HR), muscle tone, and breathing patterns are recorded simultaneously using sensors attached to a patient sleeping in a laboratory environment. This is followed by sleep staging, the manual categorization of time periods into different sleep stages based on a trained sleep technician’s visual inspection of the collected physiological data. The exact sleep staging methodology depends on the specific framework; under the American Academy of Sleep Medicine’s (AASM) [2] conventions, recordings are divided into 30-second windows, or "epochs", with electroencephalography (EEG), electromyography (EMG), and electro-oculography (EOG) used to classify epochs into three stages: awake, rapid eye movement (REM) sleep, and non-REM (NREM) sleep, which can be further subcategorized as N1, N2, and N3 sleep, with N1 as the lightest and N3 as the deepest NREM sleep. EEG is the principal signal used for sleep staging [2, 3, 4].
@@ 1.2 Wearable Devices and Automated Sleep Staging
The need for specialized personnel, equipment, and laboratory space makes PSGs expensive to conduct and difficult to scale. The cumbersome sensor setup and unfamiliar resting environment may discomfort patients, resulting in sleep that is unrepresentative of that in the home environment. Particularly sick or frail older adults may be physically unable to attend in-laboratory PSG, in addition to those living in rural or remote areas without PSG services. Furthermore, monitoring patients throughout the night and visually labelling sleep epochs is a time-consuming process that limits the throughput of sleep analyses.
In response to these challenges, wearable devices have been developed that record a subset of conventional PSG signals. Devices recording EEG have the advantage of directly measuring brain activity but the disadvantage of the discomfort associated with sensors on the head similar to PSGs. On the other hand, EEG-less devices are less obtrusive and easily set up for in-home use without a technician, but infer rather than directly measure sleep stage.
There has been considerable interest in pairing wearable devices with automated sleep stage inference in order to minimize the need for specialized technologists for labeling and hence improve scalability. Indeed, even for the analysis of conventional in-labratory PSG, automated sleep staging has been explored as a means of addressing the time-intensiveness of sleep epoch labelling. The focus on pattern recognition and loosely rule-based nature of sleep staging is ideal for machine learning approaches which have employed model architectures like convolutional neural networks (CNNs) [5], recurrent neural networks (RNNs) [6], hybrids of the previous two (e.g. convolutional recurrent neural networks (CRNN)) [7], transformers [8], and graph neural networks [9]. Deep learning models trained on PSG data, most notably the EEG signal or some combination of EEG, EMG, and EOG, have surpassed inter-scorer agreement levels, with 5-class accuracy and Cohen’s as high as 92% and 86% respectively, and can make predictions orders of magnitude faster than a human evaluator at the cost of sacrificed interpretability [10, 11].
@@ 1.3 Recurrent Neural Networks and Mamba
RNNs are a class of neural networks that takes in a sequence of vectors (e.g. a time series of pulse waves, or breathing) and produce a vector output per input, allowing them to execute a wide breadth of time-series tasks. Prevalent modern RNN architectures include long short term memory (LSTM) [12] and gated recurrent units (GRU) [13], which are known to forget very long-range dependencies and whose architecture cannot take advantage of GPU acceleration, limiting the training speed.
@@
Male/Female | 181/176 (51%/49%) | ||
Recording Duration [h] | 8.16 | 1.39 | 1.75 12.94 |
Awake [%] | 40.96 | 16.58 | 10.47 95.24 |
N1 Sleep [%] | 11.13 | 6.24 | 1.05 38.42 |
N2 Sleep [%] | 28.37 | 10.61 | 0.58 63.59 |
N3 Sleep [%] | 10.85 | 8.12 | 0.00 54.83 |
REM Sleep [%] | 8.70 | 6.08 | 0.00 29.14 |
@@
Body Mass Index (BMI) [kg ] | 27.88 | 5.59 | 16.90 50.40 |
@@ 1.4 Contribution
This paper applies Mamba-based deep learning approaches to the problem of sleep staging using data from the ANNE One [18] sensor. ANNE One, as seen in Figure 1A, is a flexible, minimally-intrusive, and clinical-grade wireless dual-module wearable system by Sibel Health composed of a finger module, which measures photoplethysmography (PPG) and temperature, and a separate module attached to the chest via an adhesive, which measures electrocardiography (ECG), triaxial accelerometry, and temperature. The raw signals of a typical recording are plotted in Figure 1B alongside its hypnogram.
A prior study centered on ANNE One [19] exists, which primarily emphasized the development of person-specific models which were trained and evaluated on data from the same patient, with a population model trained and evaluated on data from different patients showing much poorer performance. We aim to develop a superior population model to achieve accurate sleep staging from ANNE One signals recorded from previously unseen patients.
There is a considerable body of work applying machine learning to the problem of sleep stage inference from non-EEG signals. For the most part, these have been trained and tested using the cardiopulmonary sensors of conventional PSG [20, 21, 22, 23, 24, 25, 26], which are typically artifact-free. Relatively few studies have applied these approaches to data from wearable devices, which are in general more prone to poor signal quality and signal loss; consequently, such studies often need to reject epochs of poorer-quality signal prior to model training and prediction. Moreover, these works have focused on analysis of data from single sensors, typically ECG or PPG [27, 28]. Finally, models developed for wearable sensors have largely trained on data from relatively young and healthy volunteers, rather than patients presenting to sleep clinics, raising a question of generalizability to clinical populations.
The current study builds on and extends this body of work in several ways. First, we leverage the ANNE One devices’ multi-sensor nature, which may in principle allow for predictions even in settings where one or more sensors are reading low-quality signals. Second, we train and test on data from relatively older patients with a range of sleep disorders (41% were over the age of 65 and 46% had sleep apnea; Figure 1C) rather than younger and healthier patients, ensuring applicability to older adults with sleep complaints. This is particularly important as age, sleep, and neurological conditions can all plausibly alter the relationship between sleep stage and autonomic, cardiac, and pulmonary function, meaning that models trained on younger adults may generalize poorly to older adults. Finally, this work broadens the Mamba architecture to EEG-less sleep staging by exploring whether Mamba’s relative performance increases [16, 17] are also noticed in this context.
@@ 2.1 Dataset
ANNE wearable sensors were used to collect full-night recordings from 360 adults undergoing concurrent clinical PSG at a tertiary care clinical sleep laboratory in Sunnybrook Health Sciences Centre, Toronto, Canada. PSGs were recorded according to AASM practice parameters for in-labratory PSGs with a Grael PSG system (Compumedics, Victoria, Australia). The study was conducted in accordance with the latest version of the Declaration of Helsinki. All subjects provided written informed consent for the full study protocol.
Participants were first instrumented with the ANNE One, whose recordings were started ahead of the set-up and initiation of PSG recordings. In order to mimic what would occur in a home environment, the ANNE signals were not monitored in real time, and no attempt was made to correct sensor-related issues (e.g. poor skin contact) overnight. Upon completion, each PSG recording was manually labeled by one of four technologists according to AASM standards [2]. In addition, the portion of the ANNE recordings preceding the onset of PSG recording was labeled as wake based on direct observation of participants by study staff. PSG and wearable sensor data were automatically aligned by their shared ECG signals using cross-correlation analysis and verified by visual inspection.
Three participants who did not sleep at all were excluded from analysis, as were time periods prior to the sensor placement or after the sensors were taken off, leaving a total of 357 patients with 2914 hours of recording. The characteristics of the study’s participants are outlined in Figure 1C. Generally, older participants (age 65) and those with sleep apnea had poorer sleep quality, with lower total sleep time, poorer sleep efficiency, and increased N1 sleep time relative to other sleep stages (Appendix A).
The machine learning task is defined from the dataset as a form of multiclass 1D signal segmentation, wherein the ANNE One signals constitute the features, the PSG-annotated sleep stages form the labels, and class predictions are made per point in time, as demonstrated in Figure 2A.
@@ 2.2 Preprocessing
The ECG data were recorded at 512 Hz and preprocessed as follows. A wavelet transform was applied to raw ANNE ECG signals to identify temporal windows containing QRS complexes; within each window, raw ECG signals were searched for local maxima and minima to identify the precise timing of the R-peak. The wavelet-transformed ECG data were subsequently downsampled to 100 Hz for additional analyses. For each peak, morphological features (e.g. amplitude, sharpness) were extracted and hierarchical density-based spatial clustering of applications with noise (HDBSCAN) [29] served to group all identified peaks into a small number of clusters. Peaks occurring in trains with relatively stable RR intervals in the physiological range of human HR were identified as likely representing physiological QRS complexes. QRS complexes that clustered morphologically with these peaks, irrespective of regularity, were used to compute beat-to-beat heart rate.
@@ 2.3 Model
Two Mamba-based deep neural network architectures were investigated: a convolutional-recurrent neural network (CRNN) and a recurrent neural network (RNN) as summarized in Figures 2AB, 3. The CRNN is a typical sleep stage architecture [21, 11] that emulates the way human evaluators label PSGs: the convolutional component encodes physiological signals within each epoch while the recurrent component captures transitional relationships between sleep stages across epochs. The RNN architecture is a lightweight and parameter-scarce alternative to the CRNN, consisting of less parameters (580K vs. 70K). Since Mamba is designed for long-context pattern recognition, it could potentially interact with signals more effectively than the CRNN if given them directly without windowing.
The CRNN architecture has three input channels for time-series, frequency, and scalar data (see Figure 3): the first two are 3D inputs, with axes for the number of 30-second windows (mirroring the 30-second windows used in visual annotation of PSG data according to AASM criteria), the window size (i.e. 30 seconds worth of sequential data), and number of features, while the 2D scalar inputs have axes for the number of windows and number of features. The model comprises:
@@
A recurrent component, consisting of a bidirectional three-layer Mamba block that processes the embedding vectors sequentially, followed by an element-wise multi-layer perceptron (MLP) that diminishes the feature-size axis into logits for each class. The bidirectional Mamba block (Appendix B) is the concatenation of two Mamba blocks that process epochs in opposite directions, allowing for a widened receptive field to detect patterns in signals more effectively.
@@
A bidirectional three-layer Mamba block that processes windows sequentially.
@@
Our previous investigations [30] (Appendix C) have found engineered frequency domain and scalar features to yield significant performance benefits for GRU-based deep learning approaches. The engineered scalar features, which have also shown discriminatory value in other previous models [30], distribute differently depending on sleep stage, while changes to their frequency domains often designate sleep stage transitions (Appendix D). Thus, empirical selections of these features (Table 1) were introduced along with the time-series signals for both the CRNN and RNN.
@@
RNN CRNN Time-Series Frequency Scalar ECG* Heart Rate (ECG + PPG) ECG* Age PPG* Heart Rate (ECG) PPG* Sex X* Heart Rate (PPG) X* Mean, ECG SQI Y* Pulse Arrival Time (PAT) Y* Mean, PPG SQI Z* Euclidean Norm Minus One (ENMO)1 Z* Entropy, PPG Chest Temperature Z-Angle2 Z-Angle* Entropy, PPG* Limb Temperature Limb-Chest Temperature Difference Skew, PPG* Theta Scale, Heart Rate (ECG + PPG) Phi Kurtosis, ENMO Entropy, Z-Angle